# Induced Domain Walls of QCD Axion, and Gravitational Waves

### v1

@@ I Introduction
The fate of the string-wall network depends on the domain wall number, $N_{\mathrm{DW}}$, which is the number of domain walls attached to one cosmic string. If $N_{\mathrm{DW}}=1$, the network collapses soon after formation due to the tension of the domain walls. At the collapse, the network emits the QCD axions, which can account for dark matter Davis (1986) (see also Refs. Redi and Tesi (2023); Gonzalez et al. (2023); Aldabergenov et al. (2024) for a longer lifetime of domain walls due to inflationary dynamics). If $N_{\mathrm{DW}}>1$, the tension of domain walls attached to the strings balance with each other, and the network becomes stable. To avoid the overclosure of the universe, we need to introduce a potential bias, which makes the network unstable Sikivie (1982). However, such an additional potential generally shifts the potential minimum of the QCD axion and potentially spoils the PQ mechanism as a solution to the strong CP problem Barr and Seckel (1992); Kamionkowski and March-Russell (1992); Holman et al. (1992); Kawasaki et al. (2015); Chang and Cui (2023).
@@
The ALP domain walls cause a cosmological domain wall problem Zeldovich et al. (1974), so they need to decay. This can be achieved by introducing a potential bias in the heavy ALP potential. As a consequence, both domain wall networks eventually collapse while the QCD axion continues to solve the strong CP problem. The timing of the string wall network collapse depends on the magnitude of the potential bias. To preserve the success of Big Bang Nucleosynthesis (BBN), the heavy ALPs must decay to Standard Model particles at temperatures $T\gtrsim 10\,\mathrm{MeV}$ Kawasaki et al. (1999, 2000); Ichikawa et al. (2005). In addition, the collapse of the heavy ALP domain wall produces a large amount of gravitational waves, which can explain the background gravitational waves in the nHz frequency band suggested by recent observations of the pulsar timing array (PTA) Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023).111 See, for instance, Refs. Kitajima et al. (2024); Bai et al. (2023); Blasi et al. (2023); Ferreira et al. (2024a); Grüber et al. (2024) for the explanation of the PTA result using the collapse of axion domain walls. Moreover, as we will see later, the QCD axions produced by the collapse of the induced domain wall can explain all dark matter even with $f_{a}=\mathcal{O}(10^{9})$ GeV despite the pre-inflationary initial condition.
Our mechanism is not restricted to the QCD axion. It is also possible to create induced domain walls for string axions or axions with a very large decay constant, which would normally be challenging since the corresponding symmetry is not considered to be restored.222 It is possible in a certain setup where the initial position of the axion is set to be around the potential maximum Takahashi and Yin (2019) (see Appendix B). We will also discuss the cosmological implications of such induced axion domain walls for cosmic birefringence Takahashi and Yin (2021); Kitajima et al. (2022); Gonzalez et al. (2023); Kitajima et al. (2023); Ferreira et al. (2024b) (see also the experimental papers Minami and Komatsu (2020); Diego-Palazuelos et al. (2022); Eskilt (2022); Eskilt and Komatsu (2022); Eskilt et al. (2023)).
@@ II Model
where $V_{\mathrm{mix}}$ comes from the non-perturbative effects of QCD, $V_{\mathrm{DW}}$ from that of the hidden sector, and $N_{a}$, $N_{\phi}$, and $N_{\rm DW}$ are integers.
@@ III Induced domain walls
If a complex scalar field with a global U(1) symmetry acquires a nonzero VEV before or during inflation, the corresponding axion becomes nearly spatially homogeneous after inflation. On the other hand, if spontaneous symmetry breaking (SSB) occurs after inflation, the axion field takes spatially random values, and cosmic strings form in association with the SSB of the U(1) symmetry. In this scenario, when the axion potential develops and drives the axion towards its minima, domain walls emerge, attached to the cosmic strings. We assume these different initial conditions for $A$ and $\Phi$, respectively. Models that naturally realize these different initial conditions are presented in Appendix B.
@@
We see that the tension roughly follows $\sigma_{a}\propto a_{\mathrm{min}}^{2}$. Thus, in our scenario, the tension of the induced domain wall is generically smaller than in the conventional case, $\sigma_{a,\mathrm{max}}=8m_{a}f_{a}^{2}/N_{a}^{2}$. In Fig. 4, we show the dependence of $\kappa$ on the $N_{a}a_{\mathrm{min}}/f_{a}$. Note that $\kappa$ approaches ${\kappa}\simeq 1/4$ in the limit of $|N_{a}a_{\mathrm{min}}/f_{a}|\ll 1$, while it tends to ${\kappa}\simeq 2/\pi^{2}$ when $|N_{a}a_{\mathrm{min}}/f_{a}|\simeq 2\pi$. See Appendix A for more details on the configuration and tension of the induced domain wall.
@@
for all $z$. Thus, the perturbation behaves as a superposition of freely propagating plane waves, which implies that axion perturbations are neither enhanced nor suppressed when passing through the domain walls. In particular, an incoming wave propagating toward the positive $z$ direction simply passes through the domain wall, without reflection. If $|N_{a}a_{\mathrm{min}}/f_{a}|=\mathcal{O}(1)$, the mass squared of $\delta a$, $m_{a,\mathrm{eff}}^{2}(z)$, deviates from $m_{a}^{2}$ around the wall. Specifically, if $|N_{a}a_{\mathrm{min}}/f_{a}|>\pi$, the mass squared becomes negative around the domain wall. This phenomenon, however, merely increases the wavenumber and does not affect the transmission rate, as the effective potential for $\delta a$ remains continuous everywhere. Thus, we conclude that the existence and motion of the induced domain wall do not affect the time evolution of the axion excitation. [^5] This scenario contrasts with the so-called bubble misalignment mechanism Lee et al. (2024), where the axion mass varies across the bubble wall while the location of the potential minimum remains the same. We also note that this discussion does not hold with $|\delta aN_{a}/f_{a}|=\mathcal{O}(1)$.
@@ IV.2 Production of gravitational waves
Fig. 7: Gravitational wave spectrum from the domain wall collapse. The red line denotes the analytical estimate with $\sigma_{\phi}=6\times 10^{15}$ GeV3 and $T_{\mathrm{dec}}=100$ MeV assuming $\mathcal{A}=\tilde{\epsilon}_{\mathrm{GW}}=1$. The blue line denotes the numerical result for $\sigma_{\phi}=1.3\times 10^{15}$ GeV3 and $T_{\mathrm{dec}}=150$ MeV translated from Ref. Lee et al. (2024). The gray violins are from the NANOGrav 15 yr result Agazie et al. (2023). We used the temperature dependence of $g_{*}$ and $g_{*s}$ given in Ref. Saikawa and Shirai (2018).
@@ V Summary and discussion
In the main text, we focused on pre- and post-inflationary initial conditions for $A$ and $\Phi$, respectively. However, induced domain walls could appear in more general circumstances. For instance, $\phi$ domain walls formed by inflationary fluctuations could also induce QCD axion domain walls. In this scenario, the QCD axion cannot be the dominant component of dark matter due to the superhorizon-scale correlation of induced walls, which would lead to excessive isocurvature perturbations. Another possibility is that $a$ forms domain walls from inflationary fluctuations, while $\phi$ forms the string-wall network as discussed in the main text. This would result in two types of the $a$ domain walls: the usual one and the induced one with lower tension. Depending on the parameters, the usual $a$ domain walls might appear only in certain domains of $\phi$. Even if the usual QCD axion domain walls disappear, the induced ones persist as long as the $\phi$ walls exist. Since axions produced from induced domain walls do not possess isocurvature perturbations, the overall isocurvature perturbation could be suppressed depending on the parameters. We have employed various simplifying assumptions regarding mass and tension (see Eqs. (7) and (8)) and the form of the potential bias. As previously noted, a detailed study of domain wall evolution in a more general situation is warranted and is left for future work.
It is crucial to recognize that our mechanism naturally produces domain walls that are not accompanied by strings. In other words, there is no winding number and the axion is well-defined in the entire space. This feature is important because it allows the formation of domain walls even for string axions (see appendix B) or axions with extremely large decay constants. Such domain wall formation is typically challenging in these cases, as symmetry restoration is difficult or even impossible. Analogous to the QCD axion scenario, our induced domain wall mechanism can easily generate axion miniclusters even if the string axion or the axion with a large decay constant initially possesses a nearly homogeneous field value. Moreover, we emphasize that the tension of induced domain walls is generally lower than that of conventional domain walls, since the distance between adjacent minima is determined by the mixing between light and heavy axions. Consequently, for certain values of the mixing parameters, it becomes possible to suppress the abundance of light axions produced by the collapse of induced domain walls.
@@ Acknowledgments
This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 23KJ0088 (K.M.), 24K17039, (K.M.), 20H01894 (F.T.), 20H05851 (F.T. and W.Y.), 21K20364 (W.Y.), 22K14029 (W.Y.), and 22H01215 (W.Y.), Graduate Program on Physics for the Universe (J.L.) and JST SPRING Grant Number JPMJPS2114 (J.L.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
Appendix A Properties of induced domain walls
$\partial_{z}^{2}\bar{a}(z)$ $\simeq m_{a}^{2}\bar{a}(z)$ $(z<0)\ ,$ (49)
and we obtain the solution as
$\sigma_{a}\simeq\frac{m_{a}a_{\mathrm{min}}^{2}}{4}\ .$ (52)
Then, the tension of the induced domain wall becomes
Appendix B Models for initial conditions

### v2

@@ I Introduction
The fate of the string-wall network depends on the domain wall number, $N_{\mathrm{DW}}$, which is the number of domain walls attached to one cosmic string. If $N_{\mathrm{DW}}=1$, the network collapses soon after formation due to the tension of the domain walls. At the collapse, the network emits the QCD axions, which can account for dark matter Davis (1986) (see also Refs. Redi and Tesi (2023); Gonzalez et al. (2023); Aldabergenov et al. (2024) for a longer lifetime of domain walls due to inflationary dynamics). If $N_{\mathrm{DW}}>1$, the tension of domain walls attached to the strings balance with each other, and the network becomes stable. To avoid the overclosure of the universe, we need to introduce a potential bias, which makes the network unstable Sikivie (1982). However, such an additional potential generally shifts the potential minimum of the QCD axion and potentially spoils the PQ mechanism as a solution to the strong CP problem Barr and Seckel (1992); Kamionkowski and March-Russell (1992); Holman et al. (1992); Kawasaki et al. (2015); Chang and Cui (2023); Beyer and Sarkar (2023).
@@
The ALP domain walls cause a cosmological domain wall problem Zeldovich et al. (1974), so they need to decay. This can be achieved by introducing a potential bias in the heavy ALP potential. As a consequence, both domain wall networks eventually collapse while the QCD axion continues to solve the strong CP problem. The timing of the string wall network collapse depends on the magnitude of the potential bias. To preserve the success of Big Bang Nucleosynthesis (BBN), the heavy ALPs must decay to Standard Model particles at temperatures $T\gtrsim 10\,\mathrm{MeV}$ Kawasaki et al. (1999, 2000); Ichikawa et al. (2005). In addition, the collapse of the heavy ALP domain wall produces a large amount of gravitational waves, which can explain the background gravitational waves in the nHz frequency band suggested by recent observations of the pulsar timing array (PTA) Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023). [^1] See, for instance, Refs. Kitajima et al. (2024); Bai et al. (2023); Blasi et al. (2023); Gouttenoire and Vitagliano (2023); Ferreira et al. (2024a); Grüber et al. (2024) for the explanation of the PTA result using the collapse of axion domain walls. In this context, the formation of primordial black holes with masses of order $M_{\odot}$ was first noted in Ref. Kitajima et al. (2024). For a more detailed analysis, see Refs. Gouttenoire and Vitagliano (2023); Ferreira et al. (2024a). Note that the initial version of Ref. Gouttenoire and Vitagliano (2023) suggested that PBH overproduction ruled out the domain wall interpretation of the PTA signal, but this conclusion was revised in the subsequent version. Moreover, as we will see later, the QCD axions produced by the collapse of the induced domain wall can explain all dark matter even with $f_{a}=\mathcal{O}(10^{9})$ GeV despite the pre-inflationary initial condition.
Our mechanism is not restricted to the QCD axion. It is also possible to create induced domain walls for string axions or axions with a very large decay constant, which would normally be challenging since the corresponding symmetry is not considered to be restored222 It is possible in a certain setup where the initial position of the axion is set to be around the potential maximum Takahashi and Yin (2019) (see Appendix A). We will also discuss the cosmological implications of such induced axion domain walls for cosmic birefringence Takahashi and Yin (2021); Kitajima et al. (2022); Gonzalez et al. (2023); Kitajima et al. (2023); Ferreira et al. (2024b) (see also the experimental papers Minami and Komatsu (2020); Diego-Palazuelos et al. (2022); Eskilt (2022); Eskilt and Komatsu (2022); Eskilt et al. (2023)).
@@ II Model
where $V_{\mathrm{mix}}$ comes from the non-perturbative effects of QCD, $V_{\mathrm{DW}}$ from that of the hidden sector, and $N_{a}$, $N_{\phi}$, and $N_{\rm DW}$ are integers. This potential can be obtained by coupling $A$ to Peccei-Quinn fermions that only generate an anomaly for QCD and $\Phi$ to fermions that generate anomalies for both QCD and the hidden gauge group. A very simple model giving the origin of the setup and solving the quality problem for the QCD axion is provided in Appendix B.
@@ III Induced domain walls
If a complex scalar field with a global U(1) symmetry acquires a nonzero VEV before or during inflation, the corresponding axion becomes nearly spatially homogeneous after inflation. On the other hand, if spontaneous symmetry breaking (SSB) occurs after inflation, the axion field takes spatially random values, and cosmic strings form in association with the SSB of the U(1) symmetry. In this scenario, when the axion potential develops and drives the axion towards its minima, domain walls emerge, attached to the cosmic strings. We assume these different initial conditions for $A$ and $\Phi$, respectively. Models that naturally realize these different initial conditions are presented in Appendix A. The comprehensive study with more generic choices and initial conditions will be given elsewhere (see e.g. Ref. Lee et al. (2024a)).
@@
We see that the tension roughly follows $\sigma_{a}\propto a_{\mathrm{min}}^{2}$. Thus, in our scenario, the tension of the induced domain wall is generically smaller than in the conventional case, $\sigma_{a,\mathrm{max}}=8m_{a}f_{a}^{2}/N_{a}^{2}$. In Fig. 4, we show the dependence of $\kappa$ on the $N_{a}a_{\mathrm{min}}/f_{a}$. Note that $\kappa$ approaches ${\kappa}\simeq 1/4$ in the limit of $|N_{a}a_{\mathrm{min}}/f_{a}|\ll 1$, while it tends to ${\kappa}\simeq 2/\pi^{2}$ when $|N_{a}a_{\mathrm{min}}/f_{a}|\simeq 2\pi$. See Appendix C for more details on the configuration and tension of the induced domain wall.
@@
for all $z$. Thus, the perturbation behaves as a superposition of freely propagating plane waves, which implies that axion perturbations are neither enhanced nor suppressed when passing through the domain walls. In particular, an incoming wave propagating toward the positive $z$ direction simply passes through the domain wall, without reflection. If $|N_{a}a_{\mathrm{min}}/f_{a}|=\mathcal{O}(1)$, the mass squared of $\delta a$, $m_{a,\mathrm{eff}}^{2}(z)$, deviates from $m_{a}^{2}$ around the wall. Specifically, if $|N_{a}a_{\mathrm{min}}/f_{a}|>\pi$, the mass squared becomes negative around the domain wall. This phenomenon, however, merely increases the wavenumber and does not affect the transmission rate, as the effective potential for $\delta a$ remains continuous everywhere. Thus, we conclude that the existence and motion of the induced domain wall do not affect the time evolution of the axion excitation. [^5] This scenario contrasts with the so-called bubble misalignment mechanism Lee et al. (2024b), where the axion mass varies across the bubble wall while the location of the potential minimum remains the same. We also note that this discussion does not hold with $|\delta aN_{a}/f_{a}|=\mathcal{O}(1)$.
@@ IV.2 Production of gravitational waves
Fig. 7: Gravitational wave spectrum from the domain wall collapse. The red line denotes the analytical estimate with $\sigma_{\phi}=6\times 10^{15}$ GeV3 and $T_{\mathrm{dec}}=100$ MeV assuming $\mathcal{A}=\tilde{\epsilon}_{\mathrm{GW}}=1$. The blue line denotes the numerical result for $\sigma_{\phi}=1.3\times 10^{15}$ GeV3 and $T_{\mathrm{dec}}=150$ MeV translated from Ref. Lee et al. (2024b). The gray violins are from the NANOGrav 15 yr result Agazie et al. (2023). We used the temperature dependence of $g_{*}$ and $g_{*s}$ given in Ref. Saikawa and Shirai (2018).
@@ V Summary and discussion
In the main text, we focused on pre- and post-inflationary initial conditions for $A$ and $\Phi$, respectively. However, induced domain walls could appear in more general circumstances. For instance, $\phi$ domain walls formed by inflationary fluctuations could also induce QCD axion domain walls. In this scenario, the QCD axion cannot be the dominant component of dark matter due to the superhorizon-scale correlation of induced walls, which would lead to excessive isocurvature perturbations. Another possibility is that $a$ forms domain walls from inflationary fluctuations, while $\phi$ forms the string-wall network as discussed in the main text. This would result in two types of the $a$ domain walls: the usual one and the induced one with lower tension. Depending on the parameters, the usual $a$ domain walls might appear only in certain domains of $\phi$. Even if the usual QCD axion domain walls disappear, the induced ones persist as long as the $\phi$ walls exist. Since axions produced from induced domain walls do not possess isocurvature perturbations, the overall isocurvature perturbation could be suppressed depending on the parameters. We have employed various simplifying assumptions regarding mass and tension (see Eqs. (7) and (8)) and the form of the potential bias. As previously noted, a detailed study of domain wall evolution in a more general situation is warranted and is left for future work, one of which is given in Lee et al. (2024a).
It is crucial to recognize that our mechanism naturally produces domain walls that are not accompanied by strings. In other words, there is no winding number and the axion is well-defined in the entire space. This feature is important because it allows the formation of domain walls even for string axions (see appendix A) or axions with extremely large decay constants. Such domain wall formation is typically challenging in these cases, as symmetry restoration is difficult or even impossible. Analogous to the QCD axion scenario, our induced domain wall mechanism can easily generate axion miniclusters even if the string axion or the axion with a large decay constant initially possesses a nearly homogeneous field value. Moreover, we emphasize that the tension of induced domain walls is generally lower than that of conventional domain walls, since the distance between adjacent minima is determined by the mixing between light and heavy axions. Consequently, for certain values of the mixing parameters, it becomes possible to suppress the abundance of light axions produced by the collapse of induced domain walls.
@@
We also mention that the induced domain wall acts as a force potential sourced by the domain wall Kim et al. (2021). If the induced domain wall couples to the standard model fermions, such as electron, nucleon, or muon, then the force causes an effective magnetic field that induces the precisions of the spin towards the direction perpendicular to the domain wall. Thus, if the domain walls have sufficiently small tension to avoid the domain wall problem and survive until now, we can probe the existence of induced domain walls. In particular, by searching for the daily modulation of the effective magnetic field due to the Earth’s rotation in certain magnetometers, one can search for the domain wall at a faraway position.
@@ Acknowledgments
This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 23KJ0088 (K.M.), 24K17039, (K.M.), 20H01894 (F.T.), 20H05851 (F.T. and W.Y.), 21K20364 (W.Y.), 22K14029 (W.Y.), and 22H01215 (W.Y.), Incentive Research Fund for Young Researchers from Tokyo Metropolitan University (W.Y.), Graduate Program on Physics for the Universe (J.L.) and JST SPRING Grant Number JPMJPS2114 (J.L.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
Appendix A Models for initial conditions
@@
Appendix B Axiverse from non-abelian gauge symmetries
The effective theory discussed in this paper can be naturally realized if the Peccei-Quinn symmetry is an accidental symmetry arising from an ${\rm SU{(}}N)$ gauge theory, which solves the quality problem of the QCD axion Di Luzio et al. (2017); Lee and Yin (2019); Ardu et al. (2020); Yin (2020). In fact, our claim is more generic, and we will show that non-abelian gauge theory can provide an interesting axiverse scenario, where not only the axion mass but also the decay constant spread over wide ranges (cf. $\pi$ axiverse Alexander et al. (2024)). In Ref. Lee and Yin (2019), it was shown that ${\rm SU{(}}N)$ gauge theory generically contains an axion or ALP with good quality due to the spontaneous breaking of the accidental “baryon number” symmetry, either through confinement or Higgsing of the ${\rm SU{(}}N)$. This implies that if the fundamental theory contains many non-abelian gauge groups that disappear in the low-energy effective theory, we predict many axions or ALPs. Here, we further show, using a simple example, that the axiverse naturally provides pions from the confinement of the gauge group mixing with the high-quality axion. Since they originate from different symmetry breakings, we provide an axiverse with scales of not only mass but also decay constants spread over many orders of magnitude.
As a concrete example, we consider the model given in Ref. Yin (2020) (See also Ref. Ardu et al. (2020)), where we have a Higgs in a symmetric rank two tensor representation $(N\otimes N)_{\rm sym}$,
$H:\left(\frac{N^{2}+N}{2},1\right)$ (56)
and fermions in the representation  [^12] We can easily include the ${\rm SU{(}}2)_{L}$ and $U(1)_{Y}$ gauge field contribution.
$Q_{L}:(N,3),~{}~{}Q_{R}:(N,\bar{3}),~{}~{}\Psi:6\times(\bar{N},1).$ (57)
Here, the left and right arguments of the parentheses denote the representation under the hidden ${\rm SU{(}}N)$ QCD and the usual ${\rm SU{(}}3)_{c}$ QCD, respectively. Here the index $L,R$ represents the quarks in fundamental and anti-fundamental representations.
There is an accidental baryon number symmetry, under which $H$, $Q_{L,R}$, and $\Psi$ have baryon numbers of $2$, $1$, and $-1$, respecitvely. Interestingly, the ${\rm SU{(}}N)\to{\rm SO{(}}N)$ breaking can be obtained by the Higgs mechanism. The Higgs also gives mass to the fermions via the couplings such as $H^{*}Q_{L}Q_{R}$, $H\Psi\Psi$. At the same time, the accidental baryon number symmetry is spontaneously broken, giving rise to the QCD axion Ardu et al. (2020); Yin (2020).
Even after the Higgsing, assuming the Higgs-fermion couplings as well as the $SU(3)_{c}$ coupling are sufficiently small, the theory still has an approximate ${\rm SU{(}}12)$ symmetry, which should be further spontaneously broken due to the strong dynamics of ${\rm SO{(}}N).$ For certain Yukawa couplings of the Higgs field, the ${\rm SU{(}}3)_{c}$ gauge symmetry should not be spontaneously broken thanks to the mass persistency condition. On the other hand, the ${\rm SU{(}}12)$ includes the symmetry that is anomalous to the ${\rm SU{(}}3)_{c}$ subgroup. If ${\rm SU{(}}12)$ is sufficiently spontaneously broken, we have a pion coupled to the Chern-Simons term of the ${\rm SU{(}}3)_{c}$ similar to the $\pi^{0}$ coupled to the QED in the Standard Model.
In the above setup, the QCD axion arises from the spontaneous breaking of the baryon number symmetry which has a good quality. On the other hand, a neutral pion can couple to the ${\rm SU{(}}3)_{c}$ Chern-Simons term, and so it behaves as an ALP that is mixed with the QCD axion. The ALP has a decay constant around the confinement scale, but the QCD axion has a decay constant around the Higgs VEV scale. Although the details of the chiral symmetry phase transition are non-trivial and require further investigation, we have shown that the Lagrangian discussed in the main part can naturally appear as the low energy effective theory from a UV completion that solves the quality problem of the QCD axion.
This is also a concrete example that if the fundamental theory contains many non-abelian gauge theories, we expect to have a much richer ALP and axion spectra and decay constants, giving rise to a new type of axiverse.
Appendix C Properties of induced domain walls
$\partial_{z}^{2}\bar{a}(z)$ $\simeq m_{a}^{2}\bar{a}(z)$ $(z<0)\ ,$ (58)
and we obtain the solution as
$\sigma_{a}\simeq\frac{m_{a}a_{\mathrm{min}}^{2}}{4}\ .$ (61)
Then, the tension of the induced domain wall becomes